# On the convergence of noisy Bayesian Optimization with Expected Improvement

### v1

@@ 3.2 Characterizations of GP prior objective
The following lemma is on the cumulative probability distribution of $I_{t}(\cdot)$ under the GP prior assumption.
@@
Under Assumption 2.4, the cumulative probability distribution of $I_{t}$ satisfies
(25) $\mathbb{P}\{I_{t}({\bm{x}})\leq a\}=\begin{cases}0,\ &a<0,\\ \Phi\left(\frac{a}{\sigma_{t}({\bm{x}})}-z_{t}({\bm{x}})\right),\ &a\geq 0.\end{cases}$
@@
Under Assumption 2.4, at a given $t$, $f({\bm{x}})\sim\mathcal{N}(\mu_{t}({\bm{x}}),\sigma_{t}({\bm{x}}))$. Since $I_{t}({\bm{x}})\geq 0$ for all ${\bm{x}}$, (25) follows immediately if $a<0$. For $a\geq 0$,
(26) $\mathbb{P}\{I_{t}({\bm{x}})\leq a\}=\mathbb{P}\{y^{+}_{t}-f({\bm{x}})\leq a\}=1-\mathbb{P}\{f({\bm{x}})\leq y^{+}_{t}-a\}.$
Using basic properties of the CDF of the standard normal distribution,
(27) $1-\mathbb{P}\{f({\bm{x}})\leq y^{+}_{t}-a\}=1-\Phi\left(\frac{y_{t}^{+}-a-\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}\right)=\Phi\left(\frac{a-y_{t}^{+}+\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}\right).$
@@
(29) $\mathbb{P}\left\{f({\bm{x}})-\mu_{t}({\bm{x}})>\sqrt{\beta}\sigma_{t}({\bm{x}})\right\}=1-\Phi\left(\sqrt{\beta}\right)\leq\frac{1}{2}e^{-\frac{\beta}{2}}.$
@@
Lemma 3.12 is extended to $\forall t$ and ${\bm{x}}_{t}$ using union bound and an increasing $\beta_{t}$ in the following lemma (see [18] for proof).
@@
Given $\delta\in(0,1)$, let $\beta=\max\{1.44,2\log(\frac{c_{\alpha}}{\delta})\}$, where constant $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Under Assumption 2.4, at given ${\bm{x}}\in C$ and $t\geq\mathbb{N}$,
@@
Consider the first probability of (34). From Lemma 3.6, $EI_{t}({\bm{x}})\geq 0$ for $\forall{\bm{x}}$ and $t$. Therefore, $\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})>0$. From Lemma 3.6, Lemma 3.10, and the monotonicity of $\Phi$, we have
@@
Second, $-\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})\geq 0$. By Lemma 3.10, we have
@@
(38) $\frac{EI_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}<\phi(z_{t}({\bm{x}}))\leq\phi(0)<1\leq w.$
Then, (38) contradicts the premise of this case. Thus, we have $y^{+}_{t}-\mu_{t}({\bm{x}})\geq 0$ and $z_{t}({\bm{x}})\geq 0$. From the definition (4) and $\Phi\in(0,1)$,
@@
where $c_{2}(w)=[1-\phi(c_{1}(w))]^{2}+[c_{1}(w)]^{2}$. The last inequality in (45) again uses Lemma 3.1. Notice that $c_{2}(w)$ increases with $w$ and for $w\geq 1.2$, $c_{2}(w)>11$. Thus, $e^{-\frac{1}{2}w^{2}}>e^{-\frac{1}{2}c_{2}(w)w^{2}}$ for $w\geq 1.2$, which simplifies (45) to
(46) $\Phi\bigl{(}-w+$ $\phi(c_{1}(w)w)\bigr{)}$ $<\frac{1+\pi}{2\pi}e^{-\frac{1}{2}w^{2}}.$
Therefore, by (44) and (46), if $w\geq 1.2$,
(47) $\mathbb{P}\left\{I_{t}({\bm{x}})<-\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})\right\}<\frac{1+\pi}{2\pi}e^{-\frac{1}{2}w^{2}}.$
@@
(48) $\mathbb{P}\left\{\left|I_{t}({\bm{x}})-EI_{t}({\bm{x}})\right|>w\sigma_{t}({\bm{x}})\right\}<\frac{1+2\pi}{2\pi}e^{-\frac{1}{2}w^{2}},$
for $w\geq 1.2$. The probability in (48) monotonically decreases with $w$. Let
(49) $\delta=\frac{1+2\pi}{2\pi}e^{-\frac{1}{2}w^{2}}.$
Taking the logarithm of (49) leads to $\log(\frac{1+2\pi}{2\pi\delta})=\frac{1}{2}w^{2}$. Let $\beta=\max\{w^{2},1.2^{2}\}$, and the proof is complete.
@@
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{1}{\delta})$. Under Assumption 2.4, at given ${\bm{x}}\in C$ and $t\geq 1$,
@@
From Lemma 3.12, with probability $\geq 1-\delta$, (28) stands. If $y^{+}_{t}-f({\bm{x}})\leq 0$, then $I_{t}({\bm{x}})=0$. Since $EI_{t}({\bm{x}})\geq 0$, (50) is trivial. If $y^{+}_{t}-f({\bm{x}})>0$, by Lemma 3.12,
(51) $y^{+}_{t}-\mu_{t}({\bm{x}})$ $=y^{+}_{t}-f({\bm{x}})+f({\bm{x}})-\mu_{t}({\bm{x}})>f({\bm{x}})-\mu_{t}({\bm{x}})$
$>-\sqrt{\beta}\sigma_{t}({\bm{x}}),$
with probability greater than $\geq 1-\frac{\delta}{2}$. From Lemma 3.2, we have
(52) $\tau\left(\frac{y^{+}_{t}-\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}\right)>\tau(-\sqrt{\beta}),$
and therefore,
(53) $EI_{t}({\bm{x}})=\sigma_{t}({\bm{x}})\tau\left(\frac{y^{+}_{t}-\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}\right)>\tau(-\sqrt{\beta})\sigma_{t}({\bm{x}}),$
with probability greater than $1-\frac{\delta}{2}$. Next, we let $w=\sqrt{\beta}$ and consider the first probability in (34). Since $\sqrt{\beta}\sigma_{t}({\bm{x}})+EI_{t}({\bm{x}})>0$, we can follow the proof procedure and reach (35) so that
@@
Applying (55) to (53) with union bound leads to
@@ 4 Convergence analysis for GP-EI
In this section, we consider the asymptotic convergence of GP-EI under Assumption 2.4. Since $y^{+}_{t}$ is the actual output of GP-EI and plays an integral part in the algorithm itself (Algorithm 1), we consider the error measure $r_{t}$ given $t$ samples defined by
@@
where $f({\bm{x}}^{*})=f^{*}$. The error (57) is a natural extension of the noiseless error $f({\bm{x}}^{+}_{t})-f({\bm{x}}^{*})$ in [14], which is also called simple regret. Hence, the asymptotic convergence in the noisy case refers to the best existing observation converging (with probability $1-\delta$ for $\forall\delta\in(0,1)$) to the vicinity of $f^{*}$ with values no greater than $f^{*}$, i.e., $r_{t}\leq 0$ as $t\to\infty$.
While (57) is not guaranteed to be positive given the nature of noisy observations, $r_{t}$ is monotonically non-increasing, i.e., $r_{t+1}\leq r_{t}$. More importantly, our asymptotic convergence analysis of $r_{t}$ can be easily used to establish convergence of various error measures under additional conditions. For instance, a lower bound of (57) in terms of $\sigma_{t}({\bm{x}}_{t+1})$ can be established if a stopping criterion such as $EI_{t}({\bm{x}}_{t+1})\geq\kappa,\kappa>0$ is used, similar to [20]. Then, $r_{t}\to 0$ can be achieved if the stopping criterion is not met for $\forall t\in\mathbb{N}$. Equivalently, one can adopt a different error measure
(58) $\tilde{r}_{t}=f({\bm{x}}_{t})-f({\bm{x}}^{*})\geq 0,$
and obtain convergence with the stopping criterion. Another error measure for GP-EI is
(59) $r_{t}^{\prime}=I_{t}({\bm{x}}^{*})=\max\{y^{+}_{t}-f({\bm{x}}^{*}),0\}\geq 0,$
which measures the error at $t$ against ${\bm{x}}^{*}$ in terms of the improvement function $I_{t}$, an integral part of the GP-EI algorithm. It is clear that $r_{t}\leq 0$ as $t\to\infty$ directly leads to the convergence of (59).
Assumption 2.42.6 and 2.7 are used in this section. In most of our analysis, we present error bound in terms of $\sigma_{t}({\bm{x}}_{t+1})$, which satisfies $\sigma_{t}({\bm{x}}_{t+1})\to 0$ as $t\to\infty$, for many commonly-used kernels and GP-based algorithms [18, 14]. In the noiseless case, the salient idea is that $\sigma_{t}({\bm{x}}_{t+1})$ can be bounded by the maximum distance between existing $t$ samples and any point $\forall{\bm{x}}\in C$, leading to convergence rate results for specific kernels. To improve noiseless error bound, we adopt the kernel assumptions in [14] and provide the convergence rates in Theorem 4.9. However, the techniques do not apply to the noisy case. Instead, we prove convergence rates in the noisy case using information theory [18] for two of the most widely used kernels: the SE kernel and Matérn kernel in Theorem 4.12.
For clarity of presentation, we briefly summarize the results for this section. Theorem 4.2 establishes the asymptotic error bounds with and without noise under GP prior assumptions. We point out that the noiseless error bound is similar to that in [14] under the RKHS assumption with no noise. Theorem 4.7 takes advantage of the exploration and exploitation properties of EI to establish a better error bound than Theorem 4.2 with and without noise (under the GP prior assumption). In Theorem 4.12, the convergence rates in the noisy case using the improved error bounds from Theorem 4.7 are given for SE and Matérn kernels. In Proposition 4.14, we demonstrate the improvement of Theorem 4.7 compared to Theorem 4.2.
@@ 4.1 Asymptotic convergence in the noisy case
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{6}{\delta})$ and $c_{t}^{\sigma}=2\log(\frac{\pi^{2}t^{2}}{2\delta})$ The error bound of the EI algorithm satisfies
(60) $\mathbb{P}\left\{r_{t}\leq c_{\tau}(\beta)\left[3\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{t-3}+(\sqrt{\beta}+\phi(0))\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right]\right\}\geq 1-\delta,$
for given $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$, where $c_{\tau}(\beta)=\frac{\tau(\sqrt{\beta})}{\tau(-\sqrt{\beta})}$ and $t_{k}\in[\frac{t}{3}-1,t]$. If there is no observation noise, i.e., $\sigma=0$, let $\beta=2\log(\frac{2}{\delta})$. Then,
(61) $\mathbb{P}\left\{r_{t}\leq c_{\tau}(\beta)\left[2\frac{2M}{t-2}+(\sqrt{\beta}+\phi(0))\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right]\right\}\geq 1-\delta,$
@@
Since $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$, by Lemma 3.1, $\mathbb{P}\left\{\epsilon_{t}\leq\sqrt{w}\sigma\right\}=1-\Phi(-\sqrt{w})\geq 1-\frac{1}{2}e^{-\frac{1}{2}w}$, for any given $t\in\mathbb{N}$, where $w>0$ is a scalar. Similar to Lemma 3.14, if $w_{t}=2\log(\frac{\pi^{2}t^{2}}{6\delta})$, we can use union bound and write $\mathbb{P}\left\{|\epsilon_{t}|\leq\sqrt{w_{t}}\sigma\right\}\geq 1-\delta$, for all $t\in\mathbb{N}$. Let $c^{\sigma}_{t}=w_{t}=2\log(\frac{\pi^{2}t^{2}}{2\delta})$ so that
@@
for all $t\in\mathbb{N}$. From Lemma 4.1, (62), and union bound, with probability greater than $1-\frac{\delta}{3}$,
@@
Consider the observation noise with $\sigma>0$. Let $A_{k}$ be the index set for the first $k$ samples, where $|A_{k}|=k$, $k\in\mathbb{N}$. For a sequence of iterates ${\bm{x}}_{i},i\in A_{k}$, the probability of $f({\bm{x}}_{i})<y_{i}=f({\bm{x}}_{i})+\epsilon_{i}$ for all $i\in A_{k}$ is $(\frac{1}{2})^{k}$, thanks to the i.i.d. $\epsilon_{i}$. Equivalently, the probability that there exists $i\in A_{k}$ such that $f({\bm{x}}_{i})\geq y_{i}$ is $1-(\frac{1}{2})^{k}$. Let $k=[\frac{t}{3}]$ where $[x]$ is the largest integer smaller than $x$. Then, $3k\leq t\leq 3(k+1)$. For $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$, $(\frac{1}{2})^{k}\leq(\frac{1}{2})^{\frac{t}{3}-1}\leq\frac{\delta}{3}$. Thus, using union bound, there exists $k\leq t_{k}\leq 3k$ so that
@@
The choice of $\beta$ satisfies both Lemma 3.12 and 3.17 with $\frac{\delta}{6}$ each. Using $\frac{\delta}{6}$ and $t_{k}$ in Lemma 3.17, with probability greater than $1-\frac{\delta}{6}$,
@@
Using $\frac{\delta}{6}$ and $t_{k}$ in Lemma 3.12, with probability greater than $1-\frac{5\delta}{6}$,
@@
for $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$. Applying (66) to (65) leads to
@@
for $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$, where $\phi(\cdot)\leq\phi(0)$ is used.
@@
The constants $c_{\tau}(\beta)$ is dependent on $\delta$ and increases quickly as $\delta$ decreases. The step number $t_{k}$ increases with $t$. Therefore, if $\sigma_{t}({\bm{x}}_{t+1})\to 0$, $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\to 0$.
@@
The noiseless case in Theorem 4.2 is similar to Theorem 2 in [14], which assumes the objective in RKHS. Using the kernel assumptions (1)-(4) on $k({\bm{x}},{\bm{x}}^{\prime})$ in [14], Theorem 4.2 leads to the $k^{-\frac{\min\{\nu,1\}}{d}}\log^{\beta}(k)$ rate for $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$ where $k\leq t_{k}\leq 3k$, and the same rate for $r_{t}$ as in [14]. We formalize the rates in the noiseless case in Theorem 4.9. However, these rates do not hold in the noisy case as $\sigma_{t}$ does not follow Lemma 7 in [14] anymore. The noisy case, where additional uncertainty is introduced into GP-EI, generates a larger upper bound with a larger $t$ for the same probability. We prove convergence rates for two kernels using information theory in the noisy case in Theorem 4.12.

### v2

@@ 3.2 Characterizations of GP prior objective
The following lemma characterizes the CDF of $I_{t}(\cdot)$ under the GP prior assumption.
@@
Under Assumption 2.5, the probability distribution of $I_{t}$ satisfies
$\mathbb{P}\{I_{t}({\bm{x}})\leq a\}=\begin{cases}0,\ &a<0,\\ \Phi\left(\frac{a}{\sigma_{t}({\bm{x}})}-z_{t}({\bm{x}})\right),\ &a\geq 0.\end{cases}$
@@
Under Assumption 2.5, at a given $t$, $f({\bm{x}})\sim\mathcal{N}(\mu_{t}({\bm{x}}),\sigma_{t}({\bm{x}}))$. Since $I_{t}({\bm{x}})\geq 0$ for all ${\bm{x}}$, (3.10) follows immediately if $a<0$. For $a\geq 0$,
$\mathbb{P}\{I_{t}({\bm{x}})\leq a\}=\mathbb{P}\{y^{+}_{t}-f({\bm{x}})\leq a\}=1-\mathbb{P}\{f({\bm{x}})\leq y^{+}_{t}-a\}.$
Using basic properties of the standard normal CDF,
$1-\mathbb{P}\{f({\bm{x}})\leq y^{+}_{t}-a\}=1-\Phi\left(\frac{y_{t}^{+}-a-\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}\right)=\Phi\left(\frac{a-y_{t}^{+}+\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}\right).$
@@
(23) $\mathbb{P}\left\{f({\bm{x}})-\mu_{t}({\bm{x}})>\sqrt{\beta}\sigma_{t}({\bm{x}})\right\}\leq\frac{1}{2}e^{-\frac{\beta}{2}}.$
@@
Lemma 3.12 is can be extended to hold over all $t\in\mathbb{N}$ and ${\bm{x}}_{t}$ using union bound and an increasing $\beta_{t}$ in the following lemma (see [18] for proof).
@@
Given $\delta\in(0,1)$, let $\beta=\max\{1.44,2\log(\frac{c_{\alpha}}{\delta})\}$, where constant $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Under Assumption 2.5, at given ${\bm{x}}\in C$ and $t\in\mathbb{N}$,
@@
Consider the first probability in (28). From Lemma 3.6, $EI_{t}({\bm{x}})\geq 0$ for $\forall{\bm{x}}$ and $t$. Therefore, $\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})>0$. From Lemma 3.6, Lemma 3.10, and the monotonicity of $\Phi$, we have
@@
Second, let us consider the premise $-\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})\geq 0$. By Lemma 3.10, we have
@@
(32) $\frac{EI_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}<\phi(z_{t}({\bm{x}}))\leq\phi(0)<1\leq w,$
which contradicts the premise of this case. Thus, we have $y^{+}_{t}-\mu_{t}({\bm{x}})\geq 0$ (and $z_{t}({\bm{x}})\geq 0$). From the definition (4), since $\Phi\in(0,1)$,
@@
$\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}(w-\phi(c_{1}(w)w))^{2}}\phi(c_{1}(w)w)\leq\Phi(-w)+\frac{1}{2\pi}e^{-\frac{1}{2}((1-\phi(c_{1}(w)))w)^{2}}e^{-\frac{1}{2}(c_{1}(w)w)^{2}}$
$\leq\Phi(-w)+\frac{1}{2\pi}e^{-\frac{1}{2}c_{2}(w)w^{2}}\leq\frac{1}{2}e^{-\frac{1}{2}w^{2}}+\frac{1}{2\pi}e^{-\frac{1}{2}c_{2}(w)w^{2}},$
where $c_{2}(w)=[1-\phi(c_{1}(w))]^{2}+[c_{1}(w)]^{2}$. The last inequality in (39) again uses Lemma 3.1. Notice that $c_{2}(w)$ increases with $w$ and for $w\geq 1.2$, $c_{2}(w)>1$. Thus, $e^{-\frac{1}{2}w^{2}}>e^{-\frac{1}{2}c_{2}(w)w^{2}}$ for $w\geq 1.2$, which simplifies (39) to
(40) $\Phi\bigl{(}-w+$ $\phi(c_{1}(w)w)\bigr{)}$ $<c_{\pi 1}e^{-\frac{1}{2}w^{2}}.$
where $c_{\pi 1}=\frac{1+\pi}{2\pi}$. Therefore, by (38) and (40), if $w\geq 1.2$,
(41) $\mathbb{P}\left\{I_{t}({\bm{x}})<-\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})\right\}<c_{\pi 1}e^{-\frac{1}{2}w^{2}}.$
@@
(42) $\mathbb{P}\left\{\left|I_{t}({\bm{x}})-EI_{t}({\bm{x}})\right|>w\sigma_{t}({\bm{x}})\right\}<c_{\alpha}e^{-\frac{1}{2}w^{2}},$
where $c_{\alpha}=\frac{1+2\pi}{2\pi}$ for $w\geq 1.2$. The probability in (42) monotonically decreases with $w$. Let $\delta=c_{\alpha}e^{-\frac{1}{2}w^{2}}$. Then, taking the logarithm of $\delta$ leads to $\log(\frac{1+2\pi}{2\pi\delta})=\frac{1}{2}w^{2}$. Let $\beta=\max\{w^{2},1.2^{2}\}$, and the proof is complete.
@@
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{1}{\delta})$. Under Assumption 2.5, at given ${\bm{x}}\in C$ and $t\geq 1$, we have
@@
We consider two cases. First, if $y^{+}_{t}-f({\bm{x}})\leq 0$, then $I_{t}({\bm{x}})=0$. Since $EI_{t}({\bm{x}})\geq 0$, (43) stands with probability $1$.
Second, if $y^{+}_{t}-f({\bm{x}})>0$, then
(44) $y^{+}_{t}-\mu_{t}({\bm{x}})$ $=y^{+}_{t}-f({\bm{x}})+f({\bm{x}})-\mu_{t}({\bm{x}})>f({\bm{x}})-\mu_{t}({\bm{x}}).$
From the one-side inequality in Lemma 3.12, (44) implies
(45) $\mathbb{P}\{y^{+}_{t}-\mu_{t}({\bm{x}})>-\sqrt{\beta}\sigma_{t}({\bm{x}})\}\geq 1-\frac{\delta}{2},$
Then, from Lemma 3.2 the monotonicity of $\tau(\cdot)$, we have
(46) $\tau\left(z_{t}({\bm{x}})\right)>\tau(-\sqrt{\beta}),$
where $z_{t}({\bm{x}})=\frac{y^{+}_{t}-\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}$,with probability greater than $1-\frac{\delta}{2}$. Since $EI_{t}({\bm{x}})=\sigma_{t}({\bm{x}})\tau(z_{t}({\bm{x}}))$, we can write
(47) $EI_{t}({\bm{x}})=\sigma_{t}({\bm{x}})\tau\left(z_{t}({\bm{x}})\right)>\tau(-\sqrt{\beta})\sigma_{t}({\bm{x}}),$
with probability greater than $1-\frac{\delta}{2}$. Next, we let $w=\sqrt{\beta}$ and consider the first probability in (28). Since $\sqrt{\beta}\sigma_{t}({\bm{x}})+EI_{t}({\bm{x}})>0$, we can follow the proof of Lemma 3.15 for this case and write similarly to (29) that
@@
Applying (49) to (47) by eliminating $\sigma_{t}({\bm{x}})$ and using union bound, we have
@@ 4 Convergence analysis for GP-EI
In this section, we provide the asymptotic convergence analysis for GP-EI under Assumption 2.5. Since $y^{+}_{t}$ is the best existing observation and an integral part in EI itself (4), we use the error measure $r_{t}$, given $t$ samples, defined by
@@
where $f({\bm{x}}^{*})=f^{*}$. The error (51) is a natural extension of the noiseless error $f({\bm{x}}^{+}_{t})-f({\bm{x}}^{*})$ in [17] to the noisy case, which is also called simple regret. Hence, the asymptotic convergence in the noisy case refers to convergnece of $y^{+}_{t}$ (with probability $1-\delta$ for $\forall\delta\in(0,1)$) to the vicinity of $f^{*}$ with values no greater than $f^{*}$, i.e., $r_{t}\leq 0$ as $t\to\infty$. We note that $r_{t}$ is monotonically non-increasing, i.e., $r_{t+1}\leq r_{t}$. Our asymptotic convergence results of $r_{t}$ are general and can be used to establish convergence of other error measures, as elaborated in Remark 4.4.
Assumptions 2.52.7 and 2.8 are used in this section. In most of our analysis, we present error bounds in terms of $\sigma_{t}({\bm{x}}_{t+1})$, which satisfies $\sigma_{t}({\bm{x}}_{t+1})\to 0$ as $t\to\infty$, for many commonly-used kernels and GP-based algorithms [18, 17]. In the noiseless case, the salient idea is that $\sigma_{t}({\bm{x}}_{t+1})$ can be bounded by the distance between existing $t$ samples and any ${\bm{x}}\in C$, leading to concrete convergence rate characterizations for specific kernels. For an easy comparison, we adopt the kernel assumptions in [17] and provide the convergence rates in Theorem 4.9. However, the techniques do not apply to the noisy case. Instead, we prove convergence rates in the noisy case using information theory [18] for two of the most widely used kernels: the SE kernel and Matérn kernel in Theorem 4.12.
For clarity of presentation, we briefly summarize the results in this section. Theorem 4.2 establishes the asymptotic error bounds with and without noise under the GP prior assumption. We point out that the noiseless error bound is the same as that in [17]. Theorem 4.7 takes advantage of the exploration and exploitation properties of EI to establish a better error bound than Theorem 4.2 with and without noise (under the GP prior assumption). Theorem 4.9 gives the convergence rate in the noiseless case under the GP prior assumption. In Theorem 4.12, the convergence rates in the noisy case using the improved error bounds from Theorem 4.7 are given for SE and Matérn kernels. In Proposition 4.14, we elucidate the improvement of Theorem 4.7 compared to Theorem 4.2.
@@ 4.1 Asymptotic convergence in the noisy case
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{6}{\delta})$ and $c_{t}^{\sigma}=2\log(\frac{\pi^{2}t^{2}}{2\delta})$. The error bound of GP-EI satisfies
(52) $\mathbb{P}\left\{r_{t}\leq c_{\tau}(\beta)\left[6\frac{M+\sqrt{c_{t}^{\sigma}}\sigma}{t-3}+(\sqrt{\beta}+\phi(0))\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right]\right\}\geq 1-\delta,$
for given $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$, where $c_{\tau}(\beta)=\frac{\tau(\sqrt{\beta})}{\tau(-\sqrt{\beta})}$ and $t_{k}\in[\frac{t}{3}-1,t]$. If there is no observation noise, choose $\beta=2\log(\frac{2}{\delta})$. Then,
(53) $\mathbb{P}\left\{r_{t}\leq c_{\tau}(\beta)\left[4\frac{M}{t-2}+(\sqrt{\beta}+\phi(0))\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right]\right\}\geq 1-\delta,$
@@
Since $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$, by Lemma 3.1, $\mathbb{P}\left\{\epsilon_{t}\leq\sqrt{w}\sigma\right\}=1-\Phi(-\sqrt{w})\geq 1-\frac{1}{2}e^{-\frac{1}{2}w}$, for any given $t\in\mathbb{N}$, where $w>0$ is a scalar. Similar to Lemma 3.14, if $w_{t}=2\log(\frac{\pi^{2}t^{2}}{6\delta})$, we can use union bound and write $\mathbb{P}\left\{|\epsilon_{t}|\leq\sqrt{w_{t}}\sigma\right\}\geq 1-\delta$, for all $t\in\mathbb{N}$. Let $c^{\sigma}_{t}=w_{t}=2\log(\frac{\pi^{2}t^{2}}{2\delta})$ such that
@@
for all $t\in\mathbb{N}$. From Lemma 4.1, (54), and union bound, we have with probability greater than $1-\frac{\delta}{3}$ that
@@
Consider the observation noise with $\sigma>0$. Let $A_{k}$ be the index set of $k$ samples, where $|A_{k}|=k$, $k\in\mathbb{N}$. For a sequence of iterates ${\bm{x}}_{i},i\in A_{k}$, the probability of $f({\bm{x}}_{i})<y_{i}=f({\bm{x}}_{i})+\epsilon_{i}$ for all $i\in A_{k}$ is $(\frac{1}{2})^{k}$, since $\epsilon_{i}$ is i.i.d. Gaussian. Consequently, the probability that there exists $i\in A_{k}$ such that $f({\bm{x}}_{i})\geq y_{i}$ is $1-(\frac{1}{2})^{k}$.
Let $k=[\frac{t}{3}]$ where $[x]$ is the largest integer smaller than $x$. Then, $3k\leq t\leq 3(k+1)$. Now, consider ${\bm{x}}_{i},y_{i}$ where $i\in[k,3k]$. From the discussion above, there exists at least $k+1$ different $i$ such that $y_{i}^{+}-y_{i+1}^{+}<\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{k}$ with probability $1-\frac{\delta}{3}$. Thus, the probability that there exists $i\in[k,3k]$ among these $k+1$ iterates such that $f({\bm{x}}_{i})\geq y_{i}$ can be obtained via union bound to be greater than $1-\frac{\delta}{3}-(\frac{1}{2})^{k}$. For $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$, we know $(\frac{1}{2})^{k}\leq(\frac{1}{2})^{\frac{t}{3}-1}\leq\frac{\delta}{3}$. Thus, there exists $k\leq t_{k}\leq 3k$ such that
@@
The choice of $\beta$ satisfies both Lemma 3.12 and 3.17 with $\frac{\delta}{6}$ each. Using $\frac{\delta}{6}$ and $t_{k}$ in Lemma 3.17, one has with probability greater than $1-\frac{\delta}{6}$ that
@@
Using $\frac{\delta}{6}$ and $t_{k}$ in Lemma 3.12, from (56), we obtain
@@
for $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$, with probability greater than $1-\frac{5\delta}{6}$. Applying (58) to (57) and using $\Phi(\cdot)\in(0,1)$ and $\phi(\cdot)\leq\phi(0)$, we have
@@
for $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$.
@@
The constant $c_{\tau}(\beta)$ in Theorem 4.2 is dependent on $\delta$ and increases quickly as $\delta$ decreases. The iteration number $t_{k}$ increases with $t$. Therefore, if $\sigma_{t}({\bm{x}}_{t+1})\to 0$, then $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\to 0$ as well.
As mentioned above, the asymptotic convergence results of $r_{t}$ can be used to show convergence of other error measures. For instance, one can use the function value $f$ instead of the observation to define $\tilde{r}_{t}=f({\bm{x}}_{t})-f({\bm{x}}^{*})\geq 0.$ The convergence result of $r_{t}$ implies the asymptotic convergence of $\tilde{r}_{t}\to 0$ if one adopts a stopping criterion $EI_{t}({\bm{x}}_{t+1})\geq\kappa,\kappa>0$, similar to [20]. Another reasonable error measure for GP-EI is $r_{t}^{\prime}=I_{t}({\bm{x}}^{*})=\max\{y^{+}_{t}-f({\bm{x}}^{*}),0\}\geq 0$, which includes the improvement function $I_{t}$, an important part of EI. Again, the convergence result $r_{t}\leq 0$ as $t\to\infty$ ensures $r_{t}^{\prime}\to 0$.
@@
Using the kernel assumptions (1)-(4) on $k({\bm{x}},{\bm{x}}^{\prime})$ in [17], Theorem 4.2 leads to the $k^{-\frac{\min\{\nu,1\}}{d}}\log^{\beta}(k)$ rate for $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$ where $k\leq t_{k}\leq 3k$, and the same rate for $r_{t}$ as in [17]. We formalize the rates in the noiseless case in Theorem 4.9. However, these rates do not hold in the noisy case as $\sigma_{t}$ does not follow Lemma 7 in [17] anymore. The noisy case generates a larger upper bound with a larger $t$ for the same probability. We prove convergence rates for two kernels using information theory in the noisy case in Theorem 4.12.